# Orbital-Selective Diffuse Magnetic Fluctuations in Sr2RuO4: a Unified Theoretical Picture

### v1

@@ Acknowledgements.
The authors thank Antoine Georges, Alexander Lichtenstein, Hugo Strand and Alaska Subedi for the fruitful discussions. E.A.S. and M.C. also acknowledge the help of the CPHT computer support team.References
@@
- Suzuki, H and Wang, L and Bertinshaw, J and Strand, Hugo UR and Käser, S and Krautloher, M and Yang, Z and Wentzell, N and Parcollet, O and Jerzembeck, F and others (2023) Suzuki, H and Wang, L and Bertinshaw, J and Strand, Hugo UR and Käser, S and Krautloher, M and Yang, Z and Wentzell, N and Parcollet, O and Jerzembeck, F and others, “Distinct spin and orbital dynamics in Sr2RuO4,” Nature Communications 14, 7042 (2023).
@@
- You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie (2021) You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie , “High-sensitivity heat-capacity measurements on Sr2RuO4 under uniaxial pressure,” Proceedings of the National Academy of Sciences 118, e2020492118 (2021).
- Kanbayasi (1976) Akio Kanbayasi, “Magnetic properties of SrRuO3 Single Crystal,” J. Phys. Soc. Jpn. 41, 1876–1878 (1976).
- Duffy et al. (2000) J. A. Duffy, S. M. Hayden, Y. Maeno, Z. Mao, J. Kulda, and G. J. McIntyre, “Polarized-Neutron Scattering Study of the Cooper-Pair Moment in ,” Phys. Rev. Lett. 85, 5412–5415 (2000).
- Deguchi et al. (2004) K. Deguchi, Z. Q. Mao, H. Yaguchi, and Y. Maeno, “Gap Structure of the Spin-Triplet Superconductor Determined from the Field-Orientation Dependence of the Specific Heat,” Phys. Rev. Lett. 92, 047002 (2004).
- Nelson et al. (2004) K. D. Nelson, Z. Q. Mao, Y. Maeno, and Y. Liu, “Odd-Parity Superconductivity in Sr2RuO4,” Science 306, 1151–1154 (2004).
- Braden, M. and Sidis, Y. and Bourges, P. and Pfeuty, P. and Kulda, J. and Mao, Z. and Maeno, Y. (2002) Braden, M. and Sidis, Y. and Bourges, P. and Pfeuty, P. and Kulda, J. and Mao, Z. and Maeno, Y., “Inelastic neutron scattering study of magnetic excitations in ,” Phys. Rev. B 66, 064522 (2002).
- Kivelson et al. (2020) Steven Allan Kivelson, Andrew Chang Yuan, Brad Ramshaw, and Ronny Thomale, “A proposal for reconciling diverse experiments on the superconducting state in Sr2RuO4,” npj Quantum Materials 5, 43 (2020).
- Braden et al. (2002) M. Braden, O. Friedt, Y. Sidis, P. Bourges, M. Minakata, and Y. Maeno, “Incommensurate Magnetic Ordering in ,” Phys. Rev. Lett. 88, 197002 (2002).
- Kunkemöller, S. and Nugroho, A. A. and Sidis, Y. and Braden, M. (2014) Kunkemöller, S. and Nugroho, A. A. and Sidis, Y. and Braden, M., “Spin-density-wave ordering in Sr1.5 studied by neutron scattering,” Phys. Rev. B 89, 045119 (2014).
@@
- Friedt, O. and Steffens, P. and Braden, M. and Sidis, Y. and Nakatsuji, S. and Maeno, Y. (2004) Friedt, O. and Steffens, P. and Braden, M. and Sidis, Y. and Nakatsuji, S. and Maeno, Y., “Strongly Enhanced Magnetic Fluctuations in a Large-Mass Layered Ruthenate,” Phys. Rev. Lett. 93, 147404 (2004).
- Steffens, P. and Sidis, Y. and Link, P. and Schmalzl, K. and Nakatsuji, S. and Maeno, Y. and Braden, M. (2007) Steffens, P. and Sidis, Y. and Link, P. and Schmalzl, K. and Nakatsuji, S. and Maeno, Y. and Braden, M., “Field-Induced Paramagnons at the Metamagnetic Transition of ,” Phys. Rev. Lett. 99, 217402 (2007).
- Steffens, P. and Friedt, O. and Sidis, Y. and Link, P. and Kulda, J. and Schmalzl, K. and Nakatsuji, S. and Braden, M. (2011) Steffens, P. and Friedt, O. and Sidis, Y. and Link, P. and Kulda, J. and Schmalzl, K. and Nakatsuji, S. and Braden, M., “Magnetic excitations in the metallic single-layer ruthenates Ca2-xSrxRuO4 studied by inelastic neutron scattering,” Phys. Rev. B 83, 054429 (2011).
- Friedt, O. and Braden, M. and André, G. and Adelmann, P. and Nakatsuji, S. and Maeno, Y. (2001) Friedt, O. and Braden, M. and André, G. and Adelmann, P. and Nakatsuji, S. and Maeno, Y., “Structural and magnetic aspects of the metal-insulator transition in ,” Phys. Rev. B 63, 174432 (2001).
- Kriener, M. and Steffens, P. and Baier, J. and Schumann, O. and Zabel, T. and Lorenz, T. and Friedt, O. and Müller, R. and Gukasov, A. and Radaelli, P. G. and Reutler, P. and Revcolevschi, A. and Nakatsuji, S. and Maeno, Y. and Braden, M. (2005) Kriener, M. and Steffens, P. and Baier, J. and Schumann, O. and Zabel, T. and Lorenz, T. and Friedt, O. and Müller, R. and Gukasov, A. and Radaelli, P. G. and Reutler, P. and Revcolevschi, A. and Nakatsuji, S. and Maeno, Y. and Braden, M., “Structural Aspects of Metamagnetism in : Evidence for Field Tuning of Orbital Occupation,” Phys. Rev. Lett. 95, 267403 (2005).
@@
- Moon (2023) Chang-Youn Moon, “Effects of orbital selective dynamical correlation on the spin susceptibility and superconducting symmetries in ,” Phys. Rev. Res. 5, L022058 (2023).
- Kaba and Sénéchal (2019) S.-O. Kaba and D. Sénéchal, “Group-theoretical classification of superconducting states of strontium ruthenate,” Phys. Rev. B 100, 214507 (2019).
- Ramires and Sigrist (2019) Aline Ramires and Manfred Sigrist, “Superconducting order parameter of : A microscopic perspective,” Phys. Rev. B 100, 104501 (2019).
- Zhang et al. (2018) Li-Da Zhang, Wen Huang, Fan Yang, and Hong Yao, “Superconducting pairing in from weak to intermediate coupling,” Phys. Rev. B 97, 060510 (2018).
- Gingras et al. (2019) O. Gingras, R. Nourafkan, A.-M. S. Tremblay, and M. Côté, “Superconducting Symmetries of from First-Principles Electronic Structure,” Phys. Rev. Lett. 123, 217005 (2019).

### v2

@@ Acknowledgements.
We thank Antoine Georges, Alexander Lichtenstein, Hugo Strand and Alaska Subedi for fruitful discussions. We acknowledge support from IDRIS/GENCI under grant number 091393, and are thankful to the CPHT computer support team for their help.References
- Keimer et al. (2015) B. Keimer, S. A. Kivelson, M. R. Norman, S. Uchida, and J. Zaanen, “From quantum matter to high-temperature superconductivity in copper oxides,” Nature 518, 179–186 (2015).
- You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie (2021) You-Sheng Li and Naoki Kikugawa and Dmitry A. Sokolov and Fabian Jerzembeck and Alexandra S. Gibbs and Yoshiteru Maeno and Clifford W. Hicks and Jörg Schmalian and Michael Nicklas and Andrew P. Mackenzie , “High-sensitivity heat-capacity measurements on Sr2RuO4 under uniaxial pressure,” Proc. Natl. Acad. Sci. U.S.A. 118, e2020492118 (2021).
- Chronister et al. (2021) Aaron Chronister, Andrej Pustogow, Naoki Kikugawa, Dmitry A. Sokolov, Fabian Jerzembeck, Clifford W. Hicks, Andrew P. Mackenzie, Eric D. Bauer, and Stuart E. Brown, “Evidence for even parity unconventional superconductivity in Sr2RuO4,” Proc. Natl. Acad. Sci. U.S.A. 118, e2025313118 (2021).
- Li et al. (2022) You-Sheng Li, Markus Garst, Jörg Schmalian, Sayak Ghosh, Naoki Kikugawa, Dmitry A Sokolov, Clifford W Hicks, Fabian Jerzembeck, Matthias S Ikeda, Zhenhai Hu, et al., “Elastocaloric determination of the phase diagram of Sr2RuO4,” Nature 607, 276–280 (2022).
- Jerzembeck et al. (2022) Fabian Jerzembeck, Henrik S. Røising, Alexander Steppke, Helge Rosner, Dmitry A. Sokolov, Naoki Kikugawa, Thomas Scaffidi, Steven H. Simon, Andrew P. Mackenzie, and Clifford W. Hicks, “The superconductivity of Sr2RuO4 under c-axis uniaxial stress,” Nat. Commun. 13, 4596 (2022).
- Palle et al. (2023) Grgur Palle, Clifford Hicks, Roser Valentí, Zhenhai Hu, You-Sheng Li, Andreas Rost, Michael Nicklas, Andrew P. Mackenzie, and Jörg Schmalian, “Constraints on the superconducting state of from elastocaloric measurements,” Phys. Rev. B 108, 094516 (2023).
- Kivelson et al. (2020) Steven Allan Kivelson, Andrew Chang Yuan, Brad Ramshaw, and Ronny Thomale, “A proposal for reconciling diverse experiments on the superconducting state in Sr2RuO4,” npj Quantum Mater. 5, 43 (2020).
- Käser et al. (2022) Stefan Käser, Hugo U. R. Strand, Nils Wentzell, Antoine Georges, Olivier Parcollet, and Philipp Hansmann, “Interorbital singlet pairing in : A Hund’s superconductor,” Phys. Rev. B 105, 155101 (2022).
- Braden et al. (2002a) M. Braden, Y. Sidis, P. Bourges, P. Pfeuty, J. Kulda, Z. Mao, and Y. Maeno, “Inelastic neutron scattering study of magnetic excitations in ,” Phys. Rev. B 66, 064522 (2002a).
- Braden et al. (2002b) M. Braden, O. Friedt, Y. Sidis, P. Bourges, M. Minakata, and Y. Maeno, “Incommensurate Magnetic Ordering in ,” Phys. Rev. Lett. 88, 197002 (2002b).
- Kunkemöller et al. (2014) S. Kunkemöller, A. A. Nugroho, Y. Sidis, and M. Braden, “Spin-density-wave ordering in Sr1.5 studied by neutron scattering,” Phys. Rev. B 89, 045119 (2014).
- Grinenko et al. (2021) Vadim Grinenko, Shreenanda Ghosh, Rajib Sarkar, Jean-Christophe Orain, Artem Nikitin, Matthias Elender, Debarchan Das, Zurab Guguchia, Felix Brückner, Mark E Barber, et al., “Split superconducting and time-reversal symmetry-breaking transitions in Sr2RuO4 under stress,” Nat. Phys. 17, 748–754 (2021).
@@
- Suzuki et al. (2023) H. Suzuki, L. Wang, J. Bertinshaw, H. U. R. Strand, S. Käser, M. Krautloher, Z. Yang, N. Wentzell, O. Parcollet, F. Jerzembeck, N. Kikugawa, A. P. Mackenzie, A. Georges, P. Hansmann, H. Gretarsson, and B. Keimer, “Distinct spin and orbital dynamics in Sr2RuO4,” Nat. Commun. 14, 7042 (2023).
@@
- Wallerberger et al. (2019) Markus Wallerberger, Andreas Hausoel, Patrik Gunacker, Alexander Kowalski, Nicolaus Parragh, Florian Goth, Karsten Held, and Giorgio Sangiovanni, “w2dynamics: Local one- and two-particle quantities from dynamical mean field theory,” Comput. Phys. Commun. 235, 388–399 (2019).
@@
- Friedt et al. (2004) O. Friedt, P. Steffens, M. Braden, Y. Sidis, S. Nakatsuji, and Y. Maeno, “Strongly Enhanced Magnetic Fluctuations in a Large-Mass Layered Ruthenate,” Phys. Rev. Lett. 93, 147404 (2004).
- Steffens et al. (2007) P. Steffens, Y. Sidis, P. Link, K. Schmalzl, S. Nakatsuji, Y. Maeno, and M. Braden, “Field-Induced Paramagnons at the Metamagnetic Transition of ,” Phys. Rev. Lett. 99, 217402 (2007).
- Steffens et al. (2011) P. Steffens, O. Friedt, Y. Sidis, P. Link, J. Kulda, K. Schmalzl, S. Nakatsuji, and M. Braden, “Magnetic excitations in the metallic single-layer ruthenates Ca2-xSrxRuO4 studied by inelastic neutron scattering,” Phys. Rev. B 83, 054429 (2011).
- Friedt et al. (2001) O. Friedt, M. Braden, G. André, P. Adelmann, S. Nakatsuji, and Y. Maeno, “Structural and magnetic aspects of the metal-insulator transition in ,” Phys. Rev. B 63, 174432 (2001).
- Kriener et al. (2005) M. Kriener, P. Steffens, J. Baier, O. Schumann, T. Zabel, T. Lorenz, O. Friedt, R. Müller, A. Gukasov, P. G. Radaelli, P. Reutler, A. Revcolevschi, S. Nakatsuji, Y. Maeno, and M. Braden, “Structural Aspects of Metamagnetism in : Evidence for Field Tuning of Orbital Occupation,” Phys. Rev. Lett. 95, 267403 (2005).